OpenStudy (anonymous):

f(x)=tan^2(x^3+1)....f'(x)=?

OpenStudy (anonymous):

I believe that you need to use the Chain Rule for that one. Try this: f(x)=tan^2(x^3+1) Let u=tan(x^3+1) .: f(x)=u^2 .: f '(x)=2 * u * (derivative of u) And I'm sure you can do the rest.